OpenStudy (anonymous):

Defibrillators are used to revive a person when their heart has stopped beating. How do they work? What care must the person administering the treatment take? Give other examples of medical devices based on the interaction between the human body and electrical currents. @Abhisar

OpenStudy (abhisar):

Defibrillators work by administering an electrical shock to the heart muscles that might jolt them into spasming and returning to their own pattern of contracting and relaxing that pumps blood through the body, particularly the to the brain, keeping the victim alive.

OpenStudy (anonymous):

@Abhisar Give other examples of medical devices based on the interaction between the human body and electrical currents.

OpenStudy (abhisar):

ECG, Pacemakers, EEG

OpenStudy (abhisar):

got it ?

OpenStudy (anonymous):

why? and how?

OpenStudy (abhisar):

What why and what how ?

OpenStudy (abhisar):

U mean working of defibrillator ?

OpenStudy (anonymous):

the pacemaker sends and electrical current to the heart to keep the heart rate from dropping to low and keeps the heart at a normal heart rate. what does the eeg and ecg do?

OpenStudy (abhisar):

\(\bigstar\) EEG stands for Electro-Encephalograph. It is an electrically operated instrumenthaving 16-30 electrodes which when attached to scalp gives information about the electric waves originating from different parts of the brain. The waves originating from brain are also called as Berger waves.

OpenStudy (abhisar):

\(\bigstar\) ECG stands for electrocardiogram. This instrument records the electrical activity of the heart.

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

got it!

OpenStudy (abhisar):

Great ! \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (anonymous):

AED is just like a defibrillator as it slows the heart rate. When you think of the electrical devices and how they are used, is the electrical pulse that is carried through the different devices the same and do they pose the same type of danger? @Abhisar

OpenStudy (abhisar):

Post ut as a new question

OpenStudy (abhisar):

*it